# Differentially Private Federated Learning of Diffusion Models for Synthetic Tabular Data Generation

### v1

@@ Abstract.
The increasing demand for privacy-preserving data analytics in finance necessitates solutions for synthetic data generation that rigorously uphold privacy standards. We introduce DP-Fed-FinDiff framework, a novel integration of Differential Privacy, Federated Learning and Denoising Diffusion Probabilistic Models designed to generate high-fidelity synthetic tabular data. This framework ensures compliance with stringent privacy regulations while maintaining data utility. We demonstrate the effectiveness of DP-Fed-FinDiff on multiple real-world financial datasets, achieving significant improvements in privacy guarantees without compromising data quality. Our empirical evaluations reveal the optimal trade-offs between privacy budgets, client configurations, and federated optimization strategies. The results affirm the potential of DP-Fed-FinDiff to enable secure data sharing and robust analytics in highly regulated domains, paving the way for further advances in federated learning and privacy-preserving data synthesis.
@@
neural networks, diffusion models, federated learning, differential privacy, synthetic data generation, financial tabular data
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; currently under review; isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Neural networksccs: Computing methodologies Latent variable modelsccs: Security and privacy Privacy-preserving protocolsccs: Computing methodologies Distributed computing methodologies
1. Introduction
The rapidly evolving landscape of financial regulations has amplified the significance of data analytics. Central banks and financial institutions extensively gather microdata to guide policy decisions, assess risks, and maintain global stability. However, this detailed and sensitive data introduces significant privacy challenges. Real-world tabular data, crucial for developing complex models, often contains sensitive information, necessitating compliance with stringent data protection as enforced by the California Consumer Privacy Act [^1] https://oag.ca.gov/privacy/ccpa or the European General Data Protection Regulation [^2] https://eur-lex.europa.eu/eli/reg/2016/679/oj. Despite these safeguards, there remain concerns about deploying Artificial Intelligence (AI) models due to risks of data leakage (Bender et al., 2021; Kairouz et al., 2019) and model attacks (Fredrikson et al., 2015; Salem et al., 2018), which could expose personally identifiable information or confidential training data (Bender et al., 2021; Kairouz et al., 2019; Fredrikson et al., 2015; Salem et al., 2018).
Figure 1. The Financial Diffusion (FinDiff) model (Sattarov et al., 2023) (top) and the model trained with Differential Privacy (DP) (Dwork et al., 2006) (bottom), applied for mixed-type tabular data generation.
A promising approach to mitigate these privacy risks involves generating high-quality synthetic data. Synthetic data, derived from generative processes that replicate the inherent properties of real data, can provide valuable insights while preserving privacy. Unlike conventional methods such as anonymization, synthetic data generation aims to capture the underlying patterns of real data without directly exposing sensitive information. This approach is particularly relevant in high-stakes domains like finance, where data sharing and utilization are heavily regulated.
Synthetic data facilitates compliant data sharing, promoting collaboration among researchers, domain experts, and institutions. It also alleviates usage restrictions, enabling flexible data analysis without violating confidentiality agreements or regulatory boundaries. Generating high-fidelity synthetic tabular data is essential for regulatory-compliant data sharing and modeling rare, impactful events such as fraud (Charitou et al., 2021; Barse et al., 2003) or diseases. Real-world tabular data presents specific challenges, including mixed attribute types, implicit relationships, and distribution imbalances, which necessitate advanced modeling techniques.
@@ 1. Introduction
Figure 2. Schematic representation of the proposed DP-Fed-FinDiff model. It illustrates how each client $\omega_{i}$ independently trains a Federated Financial Diffusion (Fed-FinDiff) (Sattarov et al., 2023, 2024) model with Differential Privacy (DP) (Dwork et al., 2006). The timesteps $X_{T},\ldots,X_{1},X_{0}$ represent different stages of latent data representations in the generative reverse diffusion process. The individual model parameters $\theta_{i}^{\omega}$ are periodically aggregated on a central server to form the consolidated model $\theta^{\xi}$, which is then redistributed back to each client for the next optimization round.
Recent advancements in deep generative models have demonstrated impressive capabilities in creating diverse and realistic content across various domains, including images (Brock et al., 2018; Rombach et al., 2022), videos (Singer et al., 2022; Yu et al., 2023), audio (Wang et al., 2023), code (Li et al., 2022; Chen et al., 2022), natural language (OpenAI, 2023; Touvron et al., 2023) and tabular data (Kotelnikov et al., 2023; Sattarov et al., 2023). Notably, Denoising Diffusion Probabilistic Models (DDPMs) have shown exceptional quality and realism in synthetic image generation (Dhariwal and Nichol, 2021; Rombach et al., 2022). Training these models requires substantial computing resources and extensive data, which presents challenges when sensitive data is distributed across multiple institutions and cannot be shared due to privacy concerns. Federated Learning (FL), proposed as a solution, allows multiple devices to collaboratively train AI models under the orchestration of a central server, keeping the training data decentralized and enhancing privacy (McMahan et al., 2017; McMahan and Ramage, 2017). However, even with decentralized training, ensuring the privacy of sensitive data remains a critical issue. This is where Differential Privacy (DP) becomes essential. Differential Privacy provides a mathematical guarantee that the inclusion of a single data point in a dataset does not significantly affect the outcome of data analysis, thereby ensuring that individual data points remain confidential. By integrating DP into FL, it is possible to enhance the privacy of decentralized data. This integration ensures that the updates sent to the central server do not reveal sensitive information about individual data points.
@@
In this work, we propose a novel learning approach that integrates: (i) Differential Privacy, (ii) Federated Learning, and (iii) Denoising Diffusion Probabilistic Models. In summary, the main contributions we present are:
@@
  • The introduction of Differential Private Federated Financial Diffusion framework (DP-Fed-FinDiff) to create synthetic tabular data with privacy protection guarantees.
@@
  • The framework allows for precise quantification and adjustment of the privacy budget to suit the unique confidentiality and privacy requirements of the finance industry.
@@
  • Comprehensive empirical evaluation of DP-Fed-FinDiff using real-world financial datasets, demonstrating its effectiveness in synthesizing high-quality, privacy-compliant data.
2. Related Work
@@
Lately, diffusion models (Cao et al., 2022; Yang et al., 2022; Croitoru et al., 2023) and federated learning (Aledhari et al., 2020; Li et al., 2021; Zhang et al., 2021a) have garnered significant research interest. The following literature review focuses on federated deep generative modeling of tabular data with differential privacy.
@@
Deep Generative Models: Xu et al. (Xu et al., 2019) introduced CTGAN, a conditional generator for tabular data, addressing mixed data types to surpass previous models’ limitations. Building on GANs for oversampling, Engelmann and Lessmann (Engelmann and Lessmann, 2021) proposed a solution for class imbalances by integrating conditional Wasserstein GANs with auxiliary classifier loss. Jordon et al. (Jordon et al., 2018) formulated PATE-GAN to enhance data synthesis privacy, providing differential privacy guarantees by modifying the PATE framework. Torfi et al. (Torfi et al., 2022) presented a differentially private framework focusing on preserving synthetic healthcare data characteristics. Lin et al. introduced DoppelGANger, a GAN-based method for generating high-fidelity synthetic time series data (Lin et al., 2019), and later analyzed the privacy risks of GAN-generated samples, highlighting vulnerabilities to membership inference attacks (Lin et al., 2021). To handle diverse data types more efficiently, Zhao et al. (Zhao et al., 2021) developed CTAB-GAN, a conditional table GAN that efficiently addresses data imbalance and distributions. Zhang et al. (Zhang et al., 2021b) offered GANBLR for a deeper understanding of feature importance, and Noock and Guillame-Bert (Nock and Guillame-Bert, 2022) proposed a tree-based approach as an interpretable alternative. Kotelnikov et al.(Kotelnikov et al., 2022) explored tabular data modeling using multinomial diffusion models (Hoogeboom et al., 2021) and one-hot encodings, while FinDiff (Sattarov et al., 2023), foundational for our framework, uses embeddings for encoding. Recent models have emerged utilizing diffusion models to address the challenges of modeling tabular data, such as class imbalance (Schreyer et al., 2024; Roy et al., 2024) or conditional tabular data synthesis (Liu and Liu, 2024).
@@
Federated Deep Generative Models: De Goede et al. in (de Goede, 2023) devise a federated diffusion model framework utilizing Federated Averaging (McMahan et al., 2017) and a UNet backbone algorithm to train DDPMs on the Fashion-MNIST and CelebA datasets. This approach reduces the parameter exchange during training without compromising image quality. Concurrently, Jothiraj and Mashhadi in (Jothiraj and Mashhadi, 2023) introduce Phoenix, an unconditional diffusion model that employs a UNet backbone to train DDPMs on the CIFAR-10 image database. Both studies underscore the pivotal role of federated learning techniques in advancing the domain. In the context of the mixed-type tabular data, Sattarov et al. (Sattarov et al., 2024) recently introduced FedTabDiff model that merges federated learning with diffusion models.
@@
Differentially Private Federated Deep Generative Models: The integration of Differential Privacy (DP) proposed by Dwork et al. (Dwork et al., 2006), into federated learning (FL) frameworks has gained considerable attention, particularly in enhancing the privacy of deep generative models (Abadi et al., 2016; Dockhorn et al., 2022). Fan et al. (Fan, 2020) provide a comprehensive survey of differentially private generative adversarial networks, emphasizing their potential in FL environments. Gargary and De Cristofaro (Gargary and De Cristofaro, 2024) extend this by systematically reviewing federated generative models, including those leveraging DP. Specific implementations like Chen et al. (Chen et al., 2020)’s gradient-sanitized approach for differentially private GANs, and Lomurno et al. (Lomurno et al., 2022)’s secure data exchange framework illustrate practical applications. Meanwhile, Augenstein et al. (Augenstein et al., 2019) discuss deep generative models in FL settings to maintain privacy across decentralized datasets. Additionally, Zhang et al. (Zhang et al., [n. d.]) demonstrate the use of federated differentially private GANs in detecting COVID-19 pneumonia, showcasing a critical healthcare application. In the financial domain, initial steps have been taken for applications such as financial risk modeling (Zheng et al., 2020), fraud detection (Byrd and Polychroniadou, 2020), or anomaly detection (Schreyer et al., 2022). Recently, Sattarov et al.(Sattarov et al., 2023) introduced a financial diffusion model (FinDiff) for generating mixed-type tabular data, later extended to federated settings(Sattarov et al., 2024). Balch et al. (Balch et al., 2024) introduced a hierarchy of privacy levels for generative methods in financial applications.
@@
These advancements highlight the synergy between differential privacy and federated learning in developing privacy-preserving deep generative models. To the best of our knowledge this is the first attempt using differentially private diffusion models in a federated learning setup for synthesizing financial mixed-type tabular data.
3. Differentially Private Federated Diffusion
@@
This section details our proposed DP-Fed-FinDiff model, which integrates Denoising Diffusion Probabilistic Models (DDPMs) with Federated Learning (FL) and enhances it with Differential Privacy (DP) for synthetic mixed-type tabular data generation.
@@
Gaussian Diffusion Models. The Denoising Diffusion Probabilistic Model (Sohl-Dickstein et al., 2015; Ho et al., 2020) operates as a latent variable model that incrementally perturbs data  $x_{0}\in\mathbb{R}^{d}$ with Gaussian noise  $\epsilon$ through a forward process and restores it using a reverse process. Starting from  $x_{0}$ , latent variables  $x_{1},\ldots,x_{T}$ are generated via a Markov Chain, transforming them into Gaussian noise  $x_{T}\sim\mathcal{N}(0,I)$ , defined as:
(1) $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I).$
@@
In this context,  $\beta_{t}$ represents the noise level at timestep  $t$ Sampling  $x_{t}$ from  $x_{0}$ is expressed as  $q(x_{t}|x_{0})=\mathcal{N}(x_{t};{\textstyle\sqrt{1-\hat{\beta_{t}}}}x_{0},\hat{\beta_{t}}I)$ , where $\hat{\beta_{t}}=1-\prod_{i=0}^{t}(1-\beta_{i})$ In the reverse process, the model denoises  $x_{t}$ to recover  $x_{0}$ A neural network parameterized by  $\theta$ is trained to approximate each step as $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ , where  $\mu_{\theta}$ and  $\Sigma_{\theta}$ are the estimated mean and covariance. According to Ho et al. (Ho et al., 2020), with  $\Sigma_{\theta}$ being diagonal,  $\mu_{\theta}$ is calculated as:
(2) $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\hat{\alpha_{t}}}}\epsilon_{\theta}(x_{t},t)).$
@@
Here,  $\alpha_{t}:=1-\beta_{t}$ $\hat{\alpha_{t}}:=\prod_{i=0}^{t}\alpha_{i}$ , and  $\epsilon_{\theta}(x_{t},t)$ represents the predicted noise component. Empirical evidence suggests that using a simplified mean squared error (MSE) loss yields better results compared to the variational lower bound  $\log p_{\theta}(x_{0})$ , as given by:
(3) $\mathcal{L}_{t}=\mathbb{E}_{x_{0},\epsilon,t}||\epsilon-\epsilon_{\theta}(x_{t},t)||_{2}^{2}.$
@@
We employ FinDiff (Sattarov et al., 2023) as the denoising diffusion probabilistic model designed for mixed-type tabular data modality.
@@
Federated Learning. The training of DDPMs is enhanced through Federated Learning (FL) (McMahan et al., 2017), which enables learning from data distributed across multiple clients, denoted as  $\{\omega_{i}\}^{\mathcal{C}}_{i=1}$ The overall dataset is divided into subsets,  $\mathcal{D}=\{\mathcal{D}_{i}\}^{\mathcal{C}}_{i=1}$ , each accessible by a single client  $\omega_{i}$ , with varied data distributions. We employ FedTabDiff (Sattarov et al., 2024), an extension of FinDiff (Sattarov et al., 2023), in a federated setting. A central FinDiff model  $f^{\xi}_{\theta}$ with parameters  $\theta^{\xi}$ is collaboratively learned by clients. Each client  $\omega_{i}$ retains a decentralized FinDiff model  $f^{\omega}_{\theta,i}$ and contributes to the central model’s training through synchronous updates across  $r=1,\dots,\mathcal{R}$ communication rounds. A subset of clients  $\omega_{i,r}\subseteq\{\omega_{i}\}^{\mathcal{C}}_{i=1}$ is selected each round, receiving the central model parameters  $\theta^{\xi}_{r}$ , performing $\gamma=1,\dots,\Gamma$ local optimization updates, and sending updated parameters back for aggregation. Fig. 2 illustrates the process. Federated Averaging (McMahan et al., 2017) is used to compute a weighted average of the updates, defined as:
(4) $\theta^{\xi}_{r+1}\leftarrow\frac{1}{|\mathcal{D}|}\sum_{i=1}^{\lambda}|\mathcal{D}_{i}|\;\theta^{\;\omega}_{i,r+1}\,,$
@@
where  $\lambda$ is the number of participating clients,  $\theta^{\xi}_{r}$ the central, and  $\theta^{\omega}_{i,r}$ the client model parameters,  $r$ the communication round,  $|\mathcal{D}|$ the total sample count, and  $|\mathcal{D}_{i}|\subseteq|\mathcal{D}|$ the number of samples for client  $\omega_{i}$
@@
Differential Privacy. The concept of Differential Privacy (DP) (Dwork et al., 2006) is a mathematical framework that ensures an algorithm’s output does not significantly change when a single data point in the input is modified, protecting individual data points from inference. Formally, a randomized algorithm $\mathcal{A}$ provides $(\varepsilon,\delta)$-differential privacy if for any two datasets $D$ and $D^{\prime}$ differing by one element, and for any subset of outputs $S\subseteq\text{Range}(\mathcal{A})$:
(5) $\mathbb{P}[\mathcal{A}(D)\in S]\leq e^{\varepsilon}\mathbb{P}[\mathcal{A}(D^{\prime})\in S]+\delta,$
@@
Federated Learning with Differential Privacy. In the proposed DP-Fed-FinDiff model, the parameter update process is modified to incorporate the Gaussian Mechanism (Dwork et al., 2014). For each minibatch, the gradient for each individual sample $\nabla\ell(x_{t},\theta)$ is computed and then clipped individually to a maximum norm $C$. These clipped gradients are accumulated into a single gradient tensor, and Gaussian noise $\mathcal{N}(0,\sigma^{2}I)$ is added. The parameter $\sigma$ is chosen based on the desired privacy budget $\varepsilon$ and $\delta$. Each client’s local differentially private model update is computed as follows:
(6) $\theta^{\omega}_{i,r+1}=\theta^{\omega}_{i,r}-\eta(\frac{1}{|B|}\sum_{x\in B}\text{clip}(\nabla\ell(x_{t},\theta^{\omega}_{i,r}),C)+\mathcal{N}(0,\sigma^{2}I)),$
@@
4. Experimental Setup
This section describes the details of the conducted experiments, encompassing used datasets, data preparation steps, model architecture including hyperparameters, and evaluation metrics.
4.1. Datasets and Data Preparation

### v2

# Federated Diffusion Modeling with Differential Privacy for Tabular Data Synthesis

@@ Abstract
The increasing demand for privacy-preserving data analytics in various domains necessitates solutions for synthetic data generation that rigorously uphold privacy standards. We introduce the DP-FedTabDiff framework, a novel integration of Differential Privacy, Federated Learning and Denoising Diffusion Probabilistic Models designed to generate high-fidelity synthetic tabular data. This framework ensures compliance with privacy regulations while maintaining data utility. We demonstrate the effectiveness of DP-FedTabDiff on multiple real-world mixed-type tabular datasets, achieving significant improvements in privacy guarantees without compromising data quality. Our empirical evaluations reveal the optimal trade-offs between privacy budgets, client configurations, and federated optimization strategies. The results affirm the potential of DP-FedTabDiff to enable secure data sharing and analytics in highly regulated domains, paving the way for further advances in federated learning and privacy-preserving data synthesis.
Index Terms:
@@
neural networks, diffusion models, federated learning, differential privacy, synthetic data generation, mixed-type tabular data
I Introduction
The increasing demand for privacy-preserving data analytics in regulated industries, especially finance, has activated interest in federated and synthetic data solutions. Financial institutions, including central banks, supreme audit institutions, and commercial entities, collect detailed microdata to inform policy, assess credit risk, and detect fraud. However, such data is inherently sensitive and subject to strict privacy regulations such as the California Consumer Privacy Act (CCPA) [^1] https://oag.ca.gov/privacy/ccpa or the European General Data Protection Regulation (GDPR) [^2] https://eur-lex.europa.eu/eli/reg/2016/679/oj.
Despite these safeguards, the use of artificial intelligence (AI) and machine learning models in sensitive domains like finance or healthcare continues to be associated with significant privacy risks. Attacks such as membership inference or model inversion [1, 2] can expose personally identifiable or proprietary information [3, 4, 1, 2]. Moreover, data centralization is often prohibited or impractical due to regulatory, legal, or operational constraints.
Figure 1: Overview of the proposed federated diffusion modeling with differential privacy (DP) for tabular data. Each entity trains a local DP diffusion model on sensitive data, and aggregation is protected by DP to mitigate model leakage. The aggregated model generates $\varepsilon$-DP compliant synthetic data for downstream tasks with formal privacy guarantees.
This makes sensitive domains such as finance and healthcare a uniquely challenging and high-impact environment for Federated Learning (FL). Here data is distributed across entities (i.e. hospitals, banks, jurisdictions), highly non-IID, often imbalanced, and governed by strict privacy obligations. Federated learning addresses these concerns by enabling decentralized model training without requiring raw data sharing [5, 6]. However, FL alone does not guarantee resistance to inference attacks, as the shared model parameters can still leak sensitive information if intercepted or analyzed by an adversary (see Figure 1, where the “devil” icons highlight such vulnerabilities). This risk motivates the integration of Differential Privacy (DP), which mathematically bounds the influence of any individual data point on the learned model [7], thereby offering an additional layer of protection and significantly strengthening the privacy guarantees of FL systems.
Complementing these privacy-preserving mechanisms, synthetic data generation has emerged as a promising strategy to support safe data sharing and experimentation. High-fidelity synthetic tabular data can preserve analytical utility while avoiding direct exposure of real records. This is especially important in finance, where generating realistic but non-identifiable data is critical for risk modeling, fraud detection [8, 9], and simulation of rare events.
@@ I Introduction
Figure 2: Schematic representation of the proposed DP-FedTabDiff model. Each client $\omega_{i}$ independently trains a Federated Tabular Diffusion (FedTabDiff) [10] model with Differential Privacy (DP) [7]. Timesteps $X_{T},\ldots,X_{0}$ represent different stages of latent data representations in the generative reverse diffusion process. Client models $\theta_{i}^{\omega}$ are periodically aggregated into a global consolidated model $\theta^{\xi}$ on a central server, which is then redistributed back to each client for the next optimization round.
Recent progress in generative modeling has shown that Denoising Diffusion Probabilistic Models (DDPMs) outperform GANs in stability, mode coverage, and sample quality, especially in complex, multimodal distributions [11, 12]. While most diffusion research has focused on image and text, recent work like FinDiff [13] has demonstrated their viability for generating high-quality financial mixed-type tabular data. Building upon this, FedTabDiff [10] extends the diffusion framework to the federated learning setting, allowing multiple institutions to train generative models collaboratively without centralizing sensitive data. However, neither approach offers formal privacy guarantees.
@@
In this work, we propose a novel learning approach that integrates: (i) Differential Privacy, (ii) Federated Learning, and (iii) Denoising Diffusion Probabilistic Models for synthesizing tabular data under regulatory constraints. Our framework is designed for environments where privacy, decentralization, and data heterogeneity are critical. The model builds upon FinDiff [13] and FedTabDiff [10] as foundational components. In summary, the main contributions we present are as follows:
@@
  • The introduction of Differentially Private Federated Tabular Diffusion framework (DP-FedTabDiff) to create synthetic tabular data with privacy protection guarantees.
@@
  • The framework allows for precise quantification and adjustment of the privacy budget to suit the unique confidentiality and privacy requirements.
@@
  • Comprehensive empirical evaluation using real-world mixed-type tabular datasets, demonstrating its effectiveness in synthesizing high-quality, privacy-compliant data.
II Related Work
@@
Lately, diffusion models [14, 15, 16] and federated learning [17, 18, 19] have garnered significant research interest. The following literature review focuses on federated deep generative modeling of tabular data with differential privacy.
@@
Deep Generative Models: Xu et al. [20] introduced CTGAN, a conditional generator for tabular data, addressing mixed data types to surpass previous models’ limitations. Building on GANs for oversampling, Engelmann and Lessmann [21] proposed a solution for class imbalances by integrating conditional Wasserstein GANs with auxiliary classifier loss. Jordon et al. [22] formulated PATE-GAN to enhance data synthesis privacy, providing DP guarantees by modifying the PATE framework. Torfi et al. [23] presented a DP framework focusing on preserving synthetic healthcare data characteristics. Lin et al. introduced DoppelGANger, a GAN-based method for generating high-fidelity synthetic time series data [24], and later analyzed the privacy risks of GAN-generated samples, highlighting vulnerabilities to membership inference attacks [25]. To handle diverse data types more efficiently, Zhao et al. [26] developed CTAB-GAN, a conditional table GAN that efficiently addresses data imbalance and distributions. Kotelnikov et al.[27] explored tabular data modeling using multinomial diffusion models [28] and one-hot encodings, while FinDiff [13], foundational for our framework, uses embeddings for encoding. Recent models have emerged utilizing diffusion models to address the challenges of modeling tabular data, such as class imbalance [29, 30].
@@
Federated Deep Generative Models: De Goede et al. in [31] introduced a federated diffusion model framework utilizing Federated Averaging [5] and a UNet backbone algorithm to train DDPMs on the Fashion-MNIST and CelebA datasets. This approach reduces the parameter exchange during training without compromising image quality. Concurrently, Jothiraj and Mashhadi in [32] introduce Phoenix, an unconditional diffusion model that employs a UNet backbone to train DDPMs on the CIFAR-10 image database. Both studies underscore the pivotal role of federated learning techniques in advancing the domain. In the context of the mixed-type tabular data, Sattarov et al. [10] recently introduced FedTabDiff model that merges federated learning with diffusion models.
@@
Differentially Private Federated Deep Generative Models: The integration of Differential Privacy (DP) proposed by Dwork et al. [7], into federated learning (FL) frameworks has gained considerable attention, particularly in enhancing the privacy of deep generative models [33, 34]. Fan et al. [35] provide a comprehensive survey of DP generative adversarial networks, emphasizing their potential in FL environments. Gargary and De Cristofaro [36] extend this by systematically reviewing federated generative models, including those leveraging DP. Specific implementations like Chen et al. [37]’s gradient-sanitized approach for DP GANs, and Lomurno et al. [38]’s secure data exchange framework illustrate practical applications. Meanwhile, Augenstein et al. [39] discuss deep generative models in FL settings to maintain privacy across decentralized datasets. Additionally, Zhang et al. [40] demonstrate the use of federated DP GANs in detecting COVID-19 pneumonia, showcasing a critical healthcare application. In the financial domain, initial steps have been taken for applications such as fraud detection [41], or anomaly detection [42].
@@
These advances highlight the synergy between differential privacy and federated learning to develop privacy-preserving deep generative models. To the best of our knowledge, this is the first attempt using DP diffusion models in a federated learning setup for synthesizing mixed-type tabular data.
III Differentially Private Federated Diffusion
@@
This section details our proposed DP-FedTabDiff model, which integrates Denoising Diffusion Probabilistic Models (DDPMs) with Federated Learning (FL) and enhances it with Differential Privacy (DP) for tabular data generation.
@@
Gaussian Diffusion Models. The Denoising Diffusion Probabilistic Model [43, 44] operates as a latent variable model that incrementally perturbs data  $x_{0}\in\mathbb{R}^{d}$ with Gaussian noise  $\epsilon$ through a forward process and restores it using a reverse process. Starting from  $x_{0}$ , latent variables  $x_{1},\ldots,x_{T}$ are generated via a Markov Chain, transforming them into Gaussian noise  $x_{T}\sim\mathcal{N}(0,I)$ , defined as:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I).$ (1)
@@
In this context,  $\beta_{t}$ represents the noise level at timestep  $t$ . Sampling  $x_{t}$ from  $x_{0}$ is expressed as $q(x_{t}|x_{0})=\mathcal{N}(x_{t};{\textstyle\sqrt{1-\hat{\beta_{t}}}}x_{0},\hat{\beta_{t}}I)$ , where $\hat{\beta_{t}}=1-\prod_{i=0}^{t}(1-\beta_{i})$ . In the reverse process, the model denoises  $x_{t}$ to recover  $x_{0}$ . A neural network parameterized by  $\theta$ is trained to approximate each step as $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ , where  $\mu_{\theta}$ and  $\Sigma_{\theta}$ are the estimated mean and covariance. According to Ho et al. [44], with  $\Sigma_{\theta}$ being diagonal,  $\mu_{\theta}$ is calculated as:
$\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\hat{\alpha_{t}}}}\epsilon_{\theta}(x_{t},t)).$ (2)
@@
Here,  $\alpha_{t}:=1-\beta_{t}$ $\hat{\alpha_{t}}:=\prod_{i=0}^{t}\alpha_{i}$ , and  $\epsilon_{\theta}(x_{t},t)$ represents the predicted noise component. Empirical evidence suggests that using a simplified MSE loss yields better results compared to the variational lower bound  $\log p_{\theta}(x_{0})$ , as given by:
[…]
@@
We employ FinDiff [13] as the denoising diffusion probabilistic model designed for mixed-type tabular data modality.
@@
Federated Learning. The training of DDPMs is enhanced through Federated Learning (FL) [5], which enables learning from data distributed across multiple clients, denoted as  $\{\omega_{i}\}^{\mathcal{C}}_{i=1}$ . The overall dataset is divided into subsets,  $\mathcal{D}=\{\mathcal{D}_{i}\}^{\mathcal{C}}_{i=1}$ , each accessible by a single client  $\omega_{i}$ , with varied data distributions. We employ FedTabDiff [10], an extension of FinDiff [13], in a federated setting. A central FinDiff model  $f^{\xi}_{\theta}$ with parameters  $\theta^{\xi}$ is collaboratively learned by clients. Each client  $\omega_{i}$ retains a decentralized FinDiff model  $f^{\omega}_{\theta,i}$ and contributes to the central model’s training through synchronous updates across  $r=1,\dots,\mathcal{R}$ communication rounds. A subset of clients  $\omega_{i,r}\subseteq\{\omega_{i}\}^{\mathcal{C}}_{i=1}$ is selected each round, receiving the central model parameters  $\theta^{\xi}_{r}$ , performing $\gamma=1,\dots,\Gamma$ local optimization updates, and sending updated parameters back for aggregation. Fig. 2 illustrates the entire process using four clients. A weighted average of the updates (i.e. Federated Averaging [5]) is used to compute the central model parameters, defined as:
$\theta^{\xi}_{r+1}\leftarrow\frac{1}{|\mathcal{D}|}\sum_{i=1}^{\lambda}|\mathcal{D}_{i}|\;\theta^{\;\omega}_{i,r+1}\,$, (4)
@@
where  $\lambda$ is the number of participating clients,  $\theta^{\xi}_{r}$ the central, and  $\theta^{\omega}_{i,r}$ the client model parameters,  $r$ the communication round,  $|\mathcal{D}|$ the total sample count, and  $|\mathcal{D}_{i}|\subseteq|\mathcal{D}|$ the number of samples for client  $\omega_{i}$ .
@@
Differential Privacy. The concept of Differential Privacy (DP) [7] is a mathematical framework that ensures an algorithm’s output does not significantly change when a single data point in the input is modified, protecting individual data points from inference. Formally, a randomized algorithm $\mathcal{A}$ provides $(\varepsilon,\delta)$-DP if for any two datasets $D$ and $D^{\prime}$ differing by one element, and for any subset of outputs $S\subseteq\text{Range}(\mathcal{A})$:
$\mathbb{P}[\mathcal{A}(D)\in S]\leq e^{\varepsilon}\mathbb{P}[\mathcal{A}(D^{\prime})\in S]+\delta,$ (5)
@@
Federated Learning with Differential Privacy. In the proposed DP-FedTabDiff model, the parameter update process is modified to incorporate the Gaussian Mechanism [45]. For each minibatch, the gradient for each individual sample $\nabla\ell(x_{t},\theta)$ is computed and then clipped individually to a maximum norm $C$. These clipped gradients are accumulated into a single gradient tensor, and Gaussian noise $\mathcal{N}(0,\sigma^{2}I)$ is added. The parameter $\sigma$ is chosen based on the desired privacy budget $\varepsilon$ and $\delta$. Each client’s local DP model update is computed as follows:
$\theta^{\omega}_{i,r+1}=\theta^{\omega}_{i,r}-\eta(\frac{1}{|B|}\sum_{x\in B}\text{clip}(\nabla\ell(x_{t},\theta^{\omega}_{i,r}),C)+\mathcal{N}(0,\sigma^{2}I)),$ (6)
@@
IV Experimental Setup
This section describes the details of the conducted experiments, encompassing datasets, data preparation steps, model architecture with hyperparameters, and evaluation metrics.
IV-A Datasets and Data Preparation